John is retired but considering the purchase of a small business. The asking price is $500,000. Scenario 1: John's bank has agreed to loan him the money at an interest rate of 7%. Scenario 2: John also has $250,000 in an investment account earning a 8.5% yield and can use it to reduce the amount of the loan. John buys the business, an the company generates $1,000,000 in revenues and $900,000 in costs (other than interest expenses). What are John's accounting profits under both Scenarios a & b?
Answer #1

In scenario 1, he has to pay the interest on $500,000 and in scenario 2 he has to pay interest on $500,000 - $250,000 = $250,000.

Scenario 1.

Interest to be paid = 7% of $500,000 = 7/100 * $500,000 = $35,000

Other costs = $900,000

So, accounting profit = revenue - total explicit costs = $1000,000 - ($900,000 + $35,000) = $65,000

Secnario 2:

Interest to be paid = 7% of $250,000 = 7/100 * $250,000 = $17,500

Other costs = $900,000

So, accounting profit = revenue - total explicit costs = $1000,000 - ($900,000 + $17,500) = $82,500

Note: Since we are calculating accounting profit we ignored opportunity costs.